Vitamin A in analysis shows that carbon=83.86%,hydrogen=10.56%and oxygen=5.58%.what is the simple formular of vitamin A


1
Expert's answer
2021-03-23T04:22:15-0400

Each percentage represents the number of grams of an element per 100 g of that compound. Convert these masses to moles using molar masses.

"\frac{83.86}{12.0107} = 6.98 \; mol \;C \\

\frac{10.56}{1.0079} = 10.47 \;mol \;H \\

\frac{5.58}{15.9994} = 0.348 \;mol \;O"

Determine the number of moles of each element related to the other elements.

"\frac{6.98}{0.348} = 20 \;C \\

\frac{10.47}{0.348} = 30 \;H \\

\frac{0.348}{0.348} = 1 \;O"

Therefore, the simple formula of vitamin A is C20H30O